August 2008

You are currently browsing the monthly archive for August 2008.

Starting today we have contractors pulling new fibre cable along our trench which provides ResNet access to the six halls of residence in Stoke Bishop (Hiatt Baker, UH, Durdham, Wills, Churchill & Badock). We’re doing this now because we’ve discovered some damage to the current fibre, and we want to get the problem fixed now (when there are very few people in residence) before the start of term.

The trench isn’t is good condition, and there is some risk that in the process of pulling new fibre the existing fibre could be broken. If so this could interrupt ResNet, phones, and/or the halls admin network for the six residences. If this does happen the contractors have equipment on standby to repair it (this could take a few hours).

In short – it is unlikely that anyone will notice this maintenance work. With a bit of luck we won’t get any breaks. Just consider the network for those six residences ‘at risk’ for the moment.

It looks like the University single-signon system isn’t working at the moment. I’m told that people are looking into it, but as far as I can tell it’s a certificate problem.

This means that the following ResNet services are unavailable as there is no way to log in at present.

  • My ResNet
  • The in room registration system
  • All our administration tools

I’ll update this post when the problem has been resolved.

Update: 2008-08-19 – 16:47
Problem is now resolved.  I don’t have many details as yet, but access to all the above tools has been restored.

There is some central database maintenance going on today, which means the following ResNet services are unavailable:

If your ResNet connection is working, it will continue to work unaffected. If you need to register your connection, move rooms, or connect a new computer to ResNet, you’ll need to wait until later this evening.

Other systems around the university may be affected by this maintenance period. For more information, see

Update: 2008-08-14 10:43
It seems that there were problems encountered during the upgrade which meant it took longer than expected.  Normal service was resumed at about 10am today (Thursday) and after a few minor wobbles we appear to be back in business.